1) True or False: For any two events, A and B, conditional probability is always less than unconditional probability, i.e. P[B|A] < P[B].
Answer #1

Clearly this is false statement

as when A and B are independent events

P(A and B) = P(A)*P(B)

P(B|A) = P(A and B)/P(A) = P(B)

hence

P(B|A) = P(B)

but there is less than sign in statement which is not always true.

Please rate